OpenStudy (anonymous):

A machinist creates a washer by drilling a hole through the center of a circular piece of metal. If the piece of metal has a radius of x+8 and the hole has a radius of x+2, what is the area of the washer?

OpenStudy (watchmath):

area of washer = area of piece of metal - area of hole

OpenStudy (anonymous):

my options are A)12πx+60π B)12πx-60π C)x^2+12πx-60π D) x^2-12πx-60π

OpenStudy (anonymous):

where does pi come into the equation?

OpenStudy (dumbcow):

area of circle = pi*radius^2 \[A = \pi(x+8)^{2} - \pi(x+2)^{2}\]

OpenStudy (jdoe0001):

 \(\bf \textit{Area of a Circular Ring}=\pi (R^2-r^2)\\ \quad \\ R = \textit{outer radius}\\ r = \textit{inner radius}\)

OpenStudy (jdoe0001):

which is the same as dumbcow said

OpenStudy (anonymous):

I'm still having trouble working it all out...

OpenStudy (jdoe0001):

hmm look at what dumbcow typed in.... all you have to do is expand it

OpenStudy (jdoe0001):

\(\bf \textit{Area of a Circular Ring}=\pi (R^2-r^2)\\ \quad \\ R = \textit{outer radius}\\ r = \textit{inner radius}\\ \quad \\ A = \pi (R^2-r^2)\implies A = \pi [(x+8)^2-(x+2)^2]\\ \quad \\ A = \pi[(x^2+16x+64)-(x^2+4x+4)]\)

OpenStudy (jdoe0001):

cancel out like-terms and add up, and expand